# Questioning Cosmic Acceleration with DESI: The Big Stall of the Universe

### v1

@@ Abstract
Figure 2: The $1\,\sigma-5\,\sigma$ regions of background quantities $\Omega_{m}(z)$, $q(z)$, $E(z)$ and $\omega(z)$ from five individual datasets in the CPL model. As a comparison, we plot $q(z)$ on a logarithmic scale in the third row. The dashed lines denote the mean values of each quantity (red), $z=0$ (blue), $\Omega_{m}=0.5$ (black), $q=0$ (green), $E=1$ (purple), $\omega_{0}=0$ (orange) and $\omega_{0}=-1$ (pink), respectively.
Model. In the framework of general relativity Einstein:1916vd , considering a homogeneous and isotropic universe, the Friedmann equations read as $H^{2}=(8\pi G\rho)/3$ and $\ddot{a}/a=-4\pi G(\rho+3p)/3$, where $H$ is the cosmic expansion rate at a scale factor $a$ and $\rho$ and $p$ are the mean energy density and pressure of different species including baryons, dark matter (DM) and DE in the late universe. Combining two Friedmann equations, the normalized Hubble parameter $E(z)\equiv H(z)/H_{0}$ for a flat CPL universe Chevallier:2000qy ; Linder:2002et is shown as
@@ Appendix A Supplementary Material for “Questioning Cosmic Acceleration with DESI: The Big Stall of the Universe”
Figure 1: The $1\,\sigma-5\,\sigma$ regions of background quantities $\Omega_{m}(z)$, $q(z)$, $E(z)$ and $\omega(z)$ from different data combinations in the CPL model. As a comparison, we plot $q(z)$ on a logarithmic scale in the third row. The dashed lines denote the mean values of each quantity (red), $z=0$ (blue), $\Omega_{m}=0.5$ (black), $q=0$ (green), $E=1$ (purple), $\omega_{0}=0$ (orange) and $\omega_{0}=-1$ (pink), respectively.

### v3

@@ Abstract
Figure 2: The $1\,\sigma-5\,\sigma$ regions of background quantities $\Omega_{m}(z)$, $q(z)$, $E(z)$ and $\omega(z)$ from five individual datasets in the CPL model. As a comparison, we plot $q(z)$ on a logarithmic scale in the third row. The dashed lines denote the mean values of each quantity (red), $z=0$ (blue), $\Omega_{m}=0.5$ (black), $q=0$ (green), $E=1$ (purple), $\omega_{0}=0$ (yellow), $\omega_{0}=0$ (orange) and $\omega_{0}=-1$ (pink), respectively.
Model. In the framework of general relativity Einstein:1916vd , considering a homogeneous and isotropic universe, the Friedmann equations read as $H^{2}=(8\pi G\rho)/3$ and $\ddot{a}/a=-4\pi G(\rho+3p)/3$, where $H$ is the cosmic expansion rate at a scale factor $a$ and $\rho$ and $p$ are the mean energy density and pressure of different species including baryons, dark matter (DM) and DE in the late universe. Combining two Friedmann equations, the normalized Hubble parameter $E(a)\equiv H(a)/H_{0}$ for a flat CPL universe Chevallier:2000qy ; Linder:2002et is shown as
@@
It is worth noting that all the datasets prefer $\omega(a)>1$ in the future, regardless of which dataset is used. This indicates that DE will behave like a “super-stiff” fluid, where the pressure is greater than its energy density. When $\omega(a)>1$ (see Fig.2), the sound speed of DE would be greater than the speed of light $c$. However, this contradicts the principle of special relativity Einstein:1905vqn , which states that no information or energy can travel faster than the speed of light.
@@ Appendix A Supplementary Material for “Questioning Cosmic Acceleration with DESI: The Big Stall of the Universe”
Figure 1: The $1\,\sigma-5\,\sigma$ regions of background quantities $\Omega_{m}(z)$, $q(z)$, $E(z)$ and $\omega(z)$ from different data combinations in the CPL model. As a comparison, we plot $q(z)$ on a logarithmic scale in the third row. The dashed lines denote the mean values of each quantity (red), $z=0$ (blue), $\Omega_{m}=0.5$ (black), $q=0$ (green), $E=1$ (purple), $\omega_{0}=0$ (yellow), $\omega_{0}=0$ (orange) and $\omega_{0}=-1$ (pink), respectively.